The formula A=P(1+r)^2 gives the amount A in dollars that P will grow to in 2 years at interest rate r (where r is given as a decimal), using compound interest. What interest rate will cause $2000 to grow to $2163.20 in 2 years?
The interest rate is___%?
(type an integer or decimal, Round to the nearest tenth of a percent if needed)

Answer by ewatrrr(24785)

 
Hi,
What interest rate will cause $2000 to grow to $2163.20 in 2 years?
2163.20/2000 = (1 + r)^2
1.0816 = (1+r)^2 |taking the square root of both sides of the EQ
1.04 = (1+ r)
.04 = r or r = 4% interest